# Principal stratification with recurrent events truncated by a terminal event: A nested Bayesian nonparametric approach

### v2

@@ 2.1 Principal stratification in continuous time
Principal stratification (Frangakis2002) is a general framework for addressing intermediate outcomes and is particularly attractive for applications with death truncation. This framework focuses on the always-survivor subgroup defined by the combination of observed and missing potential outcomes for the terminal event. Suppose the time to the first event is of interest, an existing estimand is the survivor average causal effect (SACE) for a fixed time point $t$: $\mathrm{Pr}(T_{i1}^{1}<t\mid t<D_{i}^{0},t<D_{i}^{1})-\mathrm{Pr}(T_{i1}^{0}<t\mid t<D_{i}^{0},t<D_{i}^{1})$. This quantity allows us to circumvent the potential bias by focusing on the subpopulation of individuals who would always survive regardless of the assigned treatment, i.e., both the first and second probability statements are conditioned on the same cohort of individuals. For example, Lyu2023 adopted this definition of the always-survivor and defined the SACE for recurrent events with a terminal event. In particular, they considered the causal contrast of the number of recurrent events defined as $\mathbb{E}[N_{i}^{1}(t)-N_{i}^{0}(t)\mid t<D_{i}^{0},t<D_{i}^{1}]$.
@@
Comment2019 argued that estimating the SACE at a single time point $t$ may present certain challenges. This is because it describes causal effects only for the always-survivor subpopulation at that specific moment, which they call a snapshot effect. This static approach fails to account for the sensitivity to the choice of $t$ because the focused subpopulation changes and diminishes over time, and thus overlooks how treatment effects accumulate or vary over time for a fixed subpopulation. Consequently, these snapshot estimands cannot capture the full temporal nature of treatment effects and fail to provide meaningful insights about the causal effects. Recent studies (e.g., Comment2019; Xu2022) have addressed this issue by focusing on double-indexed estimands, where the principal strata are defined by the time index $r$, independent of the choice of $t$ that indexes the time to the events of interest. This approach introduces the time-varying SACE (TV-SACE), where the population of always-survivors depends only on the chosen time $r$, that is, $\mathrm{Pr}(T_{i1}^{1}<t\mid r<D_{i}^{0},r<D_{i}^{1})-\mathrm{Pr}(T_{i1}^{0}<t\mid r<D_{i}^{0},r<D_{i}^{1})$.
The double-index approach appears promising; however, it has potential limitations when applied to recurrent events. Specifically, the traditional definition $\mathcal{AS}_{\text{trad}}(r)=\{i:r<D_{i}^{0},r<D_{i}^{1}\}$ may unnecessarily exclude units from the always-survivor stratum when there is a large gap between the time of the last recurrent event and the cutoff time $r$. Figure 1 illustrates this phenomenon: units 1–4 are excluded from the always-survivor stratum according to the standard definition, despite surviving until the final recurrent event of interest. Consequently, the inference for the stratum may become unstable as the stratum size becomes smaller. Additionally, as $r$ increases, the subpopulation of always-survivors eventually diminishes and may even be empty. These limitations motivate us to consider alternative definitions of the always-survivor subpopulation in the context of recurrent events.
Figure 1: Illustration of different definitions of always-survivor strata in recurrent event analyses with terminal events. The traditional definition is given by $\mathcal{AS}_{\text{trad}}(r)=\{i:r<D_{i}^{0},r<D_{i}^{1}\}$. Under our definition, units 1-5 belong to the always-survivor stratum, while the traditional definition excludes units 1-4, classifying only unit 5 as an always-survivor.
Specifically, we propose a partition based on the ordering of the terminal event outcome and the non-terminal recurrent event outcome, for a fixed time point $r$ with $t<r$. We define the always-survivor stratum at time $r$ as $\mathcal{AS}(r)=\quantity{i:T^{1}_{iN_{i}^{1}(r)}<D_{i}^{1},T_{iN_{i}^{0}(r)}^{0}<D_{i}^{0}}$, which consists of individuals for whom the $N_{i}^{z}(r)$-th recurrent event occurs before the terminal event under both treatment and control. The never-survivor stratum is defined as as $\mathcal{NS}(r)=\quantity{i:T^{1}_{iN_{i}^{1}(r)}=*,T_{iN_{i}^{0}(r)}^{0}=*}$, indicating individuals for whom the $N_{i}^{z}(r)$-th recurrent event would have occurred only after the terminal event under both treatment and control, and is therefore undefined as a pre-death recurrent event. See Table 1 for the definitions of the remaining two strata. Similar ideas can be found in Nevo2022; Yu2024; Zehavi2025, but their strata are defined for non-recurrent time-to-event. Note that our definition of the always-survivor stratum circumvents the aforementioned issue that it becomes scarce or even empty as time grows by considering survivorship at the exact time when the final recurrent event occurs before a prefixed time $r$. Even when the gap between the final recurrent event and the fixed time $r$ is large, our definition of strata is robust to the choice of $r$, providing a more time-stable subgroup of survival units tailored for unique features of recurrent events. Figure 1 illustrates the difference between our approach and the conventional approach (e.g., Lyu2023) in defining always-survivors.
Table 1: Stratification based on the recurrent event and terminal event.
Name Stratum definition Always-survivor $\left(\mathcal{AS}(r)\right)$ $T^{0}_{iN_{i}^{0}(r)}<D_{i}^{0}\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ T_{iN_{i}^{1}(r)}^{1}<D_{i}^{1}$ Never-survivor $\left(\mathcal{NS}(r)\right)$ $T^{0}_{iN_{i}^{0}(r)}=*\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ T_{iN_{i}^{1}(r)}^{1}=*$ Survival-harmed $\left(\mathcal{SH}(r)\right)$ $T^{0}_{iN_{i}^{0}(r)}=*\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ T_{iN_{i}^{1}(r)}^{1}<D_{i}^{1}$ Survival-protected $\left(\mathcal{SP}(r)\right)$ $T^{0}_{iN_{i}^{0}(r)}<D_{i}^{0}\penalty 10000\ \penalty 10000\ \&\penalty 10000\ \penalty 10000\ T_{iN_{i}^{1}(r)}^{1}=*$
2.2 Principal stratum causal estimands
@@
Our interest lies in evaluating the causal effect on the recurrent event process for always-survivors in Table 1, which is relevant when $N_{i}^{z}(\cdot)$ and $T_{ij}^{z}$ can be undefined due to truncation. We first define the survivor-average number of recurrence (SANR) estimands as
@@
where $\mu^{z}\quantity(t;r)=\mathbb{E}\quantity[N_{i}^{z}(t)\mid\mathcal{AS}(r)]$ for $t<r$. The function $g(\cdot,\cdot)$ determines the scale of effect measure. For example, g(x,y)=xyg(x,y)=x-y and $g(x,y)=x/y$ correspond to causal mean difference and causal rate ratio, respectively. These class of estimands represent the causal contrast in the number of event occurrences before time $t$. The second class of estimands pertain to the event rate among always-survivors, thereby answering the question of how many recurrent events we expect per unit time alive under each treatment arm. For the principal stratum $\mathcal{AS}(r)$ we define $\tau^{z}(t;r)=\mathbb{E}\bigl[T_{iN_{i}^{z}(t)}^{\,z}\mid\mathcal{AS}(r)\bigr],$ so that the ratio $\mu^{z}(t;r)\big/\tau^{z}(t;r)$ represents the average number of events per unit exposure time for an always-survivor in arm $z$. We formalize the survivor-average event-rate (SAER) estimands as
$\mathrm{SAER}(t;r)=g\quantity{\frac{\mu^{1}(t;r)}{\tau^{1}(t;r)},\,\frac{\mu^{0}(t;r)}{\tau^{0}(t;r)}}$. (2)
$\mathrm{SAER}(t;r)<1$ indicates that, per person-year alive, the treatment reduces the expected number of events relative to control, whereas a value above one indicates the opposite. The SAER estimand quantifies how many events are caused per unit follow-up, providing a direct public-health interpretation in terms of event occurrences per person-year. A more detailed discussion on the SAER is provided in Web Appendix A.
@@ 3 Bayesian Nonparametric Causal Inference
Leveraging Bayesian methods for causal inference, we first consider the joint distribution of all potential variables and covariates $(\mathbf{N}^{0}(\cdot),\mathbf{D}^{0},\mathbf{C}^{0},\mathbf{N}^{1}(\cdot),\mathbf{D}^{1},\mathbf{C}^{1},\mathbf{X},\mathbf{Z})$. We assume that the parameters governing the distribution of $\mathbf{C}^{z}$, $\mathbf{Z}$ and $\mathbf{X}$ is a priori independent of the parameters of $\mathbf{N}^{z}(\cdot)$ and $\mathbf{D}^{z}$. Under Assumption 3 and the prior independence, the censoring variables do not affect the inference of the quantities of scientific interest (e.g., Rubin1978), thus we focus on the joint distribution of $(\mathbf{N}^{0}(\cdot),\mathbf{D}^{0},\mathbf{N}^{1}(\cdot),\mathbf{D}^{1},\mathbf{Z},\mathbf{X})$. Assuming unit exchangeability, the de Finetti’s theorem implies the existence of parameters such that
@@
$\begin{split}&p\quantity(\mathbf{N}^{0}(\cdot),\mathbf{D}^{0},\mathbf{N}^{1}(\cdot),\mathbf{D}^{1},\mathbf{Z},\mathbf{X})\\ =&\int_{\Theta}p\quantity(\boldsymbol{\theta}^{*})p\quantity(\boldsymbol{\varphi}^{*})\prod_{i=1}^{n}p\quantity(N_{i}^{0}(\cdot)D_{i}^{0},N_{i}^{1}(\cdot),D_{i}^{1}\mid\mathbf{X}_{i},\boldsymbol{\theta}^{*})p(Z_{i},\mathbf{X}_{i}\mid\boldsymbol{\varphi}^{*})d\boldsymbol{\theta}^{*}d\boldsymbol{\varphi}^{*},\end{split}$ (3)
@@
where the global parameter $\boldsymbol{\theta}^{*}$ has prior $p(\boldsymbol{\theta}^{*})$. Under the ignorable treatment assignment (Assumption 2), the assignment model does not affect the inference of the quantities of scientific interest $\boldsymbol{\theta}^{*}$ either. In what follows, we will conduct a likelihood analysis for $\boldsymbol{\theta}^{*}$, assuming that the value of $\boldsymbol{\theta}^{*}$ which governed the distribution of observable data has been drawn from a prior distribution with compact support. For identification, we further assume that a frailty $\boldsymbol{\gamma}$ is included as unknown parameters in an expanded parameter set $\boldsymbol{\theta}^{*}=(\boldsymbol{\theta},\boldsymbol{\gamma})$ for $\boldsymbol{\gamma}=(\boldsymbol{\gamma}_{1},\ldots,\boldsymbol{\gamma}_{n})^{\top}$, as formalized below.
Assumption 4 (Cross-world conditional independence).
@@
For each unit $i$, there exists a bivariate frailty $\boldsymbol{\gamma}_{i}=(\gamma_{i}^{0},\gamma_{i}^{1})$ such that the joint distribution of the potential outcomes is factorized: $p\quantity(N_{i}^{0}(\cdot),D_{i}^{0},N_{i}^{1}(\cdot),D_{i}^{1}\mid\mathbf{X}_{i},\boldsymbol{\gamma}_{i})=p\quantity(N_{i}^{0}(\cdot),D_{i}^{0}\mid\mathbf{X}_{i},\gamma_{i}^{0})p\quantity(N_{i}^{1}(\cdot),D_{i}^{1}\mid\mathbf{X}_{i},\gamma_{i}^{1})$.
@@
The frailty device has been previously considered in the semi‑competing risks literature (e.g., Stensrud2017; Comment2019; Nevo2022; Lyu2023). Here, the individual frailty $\boldsymbol{\gamma}_{i}$ governs two distinct forms of dependence: (i) cross-world dependence between treatment arms, arising from the association between $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$; and (ii) within‑world dependence between the terminal event $D_{i}^{z}$ and the recurrent event process $N_{i}^{z}(\cdot)$ under treatment $z$. A popular simplification sets $\gamma_{i}^{0}=\gamma_{i}^{1}\equiv\gamma_{i}$. Although this assumption conveniently captures within-world dependence between recurrent events and the terminal event in descriptive (non-causal) analyses (e.g., Paulon2020; Tian2024), in causal analyses it also imposes the same dependence structure between cross-worlds, thereby yielding point identification of the causal contrasts. (e.g., Comment2019; Lyu2023). Because this within‑world dependence shared across different worlds rarely holds unless all effect modifiers are measured and modeled correctly, this approach imposes strong, and often opaque, assumptions on dependence structures in causal settings.
@@
Similar to Nevo2022 who addressed the non-recurrent event setting, we retain separate frailties $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$ and introduce a sensitivity parameter $\rho=\operatorname{Corr}(\gamma_{i}^{0},\gamma_{i}^{1})$ that captures their cross‑world association. This formulation separates identifiable and non‑identifiable components: the marginal distributions of $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$, and hence the within‑world dependence, are estimable from the observed data, whereas the cross‑world correlation $\rho$ is an unidentifiable, cross-world quantity. Nevo2022 specified a correlated gamma distribution for the joint frailty, fixing the location parameters and correlation parameter $\rho$ a priori and thereby constraining the model to a relatively restrictive parametric family. We consider a sensitivity analysis framework by placing a flexible non‑parametric prior on the joint distribution of $(\gamma_{i}^{0},\gamma_{i}^{1})$ conditional on a fixed $\rho$, specified in Section 3.1.2. Sensitivity analysis is carried out by varying $\rho$ and examining how posterior estimates of the causal estimands change, as illustrated in Section 5.
Assumption 4 allows us to derive the posterior distribution of $\boldsymbol{\theta}^{*}=(\boldsymbol{\theta},\boldsymbol{\gamma})$ as
$p\quantity(\boldsymbol{\theta},\boldsymbol{\gamma}\mid\mathbf{N}^{\text{obs}}(\cdot),\mathbf{D}^{\text{obs}},\mathbf{Z},\mathbf{X})\propto p\quantity(\boldsymbol{\theta},\boldsymbol{\gamma})\prod_{i=1}^{n}p\quantity(N_{i}(\cdot),D_{i}\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\theta},\gamma_{i}^{Z_{i}}),$ (4)
where the vectors with the superscript ‘obs’ denote the observed vectors of each potential outcome. To obtain the posterior distribution of the causal estimands, we then specify priors of $\boldsymbol{\theta}$ and $\boldsymbol{\gamma}$ and the likelihood for (4) and impute each unit’s missing potential outcomes by sampling from their posterior predictive distribution, $p\quantity(N_{i}^{1-Z_{i}}(\cdot),D_{i}^{1-Z_{i}}\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\theta},\boldsymbol{\gamma}_{i}^{1-Z_{i}})$. Theorem 1 states that (1) is identifiable up to the frailty $\boldsymbol{\gamma}$.

### v3

@@ 2.2 Principal stratum causal estimands in continuous time
Principal stratification (Frangakis and Rubin, 2002) is a general framework for addressing intermediate outcomes and is particularly attractive for applications with death truncation. This framework focuses on the always-survivor subgroup defined by the combination of observed and missing potential outcomes for the terminal event. Suppose the time to the first event is of interest, an existing estimand is the survivor average causal effect (SACE) for a fixed time point $t$:
$\mathrm{Pr}(T_{i1}^{1}<t\mid t<D_{i}^{0},t<D_{i}^{1})-\mathrm{Pr}(T_{i1}^{0}<t\mid t<D_{i}^{0},t<D_{i}^{1}).$
This quantity allows us to circumvent the potential bias by focusing on the subpopulation of individuals who would always survive regardless of the assigned treatment, i.e., both the first and second probability statements are conditioned on the same cohort of individuals. For recurrent event analysis, Lyu et al. (2023) adopted this definition of the always-survivor and defined the SACE for recurrent events with a terminal event. In particular, they considered the causal contrast of the number of recurrent events defined as
$\frac{\mathbb{E}[N_{i}^{1}(t)\mid t<D_{i}^{0},t<D_{i}^{1}]}{\mathbb{E}[N_{i}^{0}(t)\mid t<D_{i}^{0},t<D_{i}^{1}]}.$
@@
Comment et al. (2025) argued that estimating the SACE at a single time point $t$ may present certain challenges. This is because it describes causal effects only for the always-survivor subpopulation at that specific moment, which they call a snapshot effect. This static approach fails to account for the sensitivity to the choice of $t$ because the focused subpopulation changes and diminishes over time, and thus overlooks how treatment effects accumulate or vary over time for a fixed subpopulation. Consequently, these snapshot estimands cannot capture the full temporal nature of treatment effects and fail to provide meaningful insights about the causal effects. Comment et al. (2025) addressed this issue by focusing on double-indexed estimands, where the principal strata are defined by the time index $r$, independent of the choice of $t$ that indexes the time to the events of interest. This approach introduces the time-varying SACE (TV-SACE), where the population of always-survivors depends only on the chosen time $r$, that is,
$\mathrm{Pr}(T_{i1}^{1}<t\mid r<D_{i}^{0},r<D_{i}^{1})-\mathrm{Pr}(T_{i1}^{0}<t\mid r<D_{i}^{0},r<D_{i}^{1}).$
Their estimands focus on the first time-to-event outcome rather than recurrent events.
We adopt the double-index approach for the recurrent event analysis and define the subpopulation of always-survivors at time $r$ as:
$\mathcal{AS}(r)=\{i:r<D_{i}^{0},r<D_{i}^{1}\}\penalty 10000\ \penalty 10000\ \text{for}\penalty 10000\ \penalty 10000\ t<r.$
@@
Our interest lies in evaluating the causal effect on the recurrent event process for those in $\mathcal{AS}(r)$, which is relevant when $N_{i}^{z}(\cdot)$ and $T_{ij}^{z}$ can be undefined due to truncation. We define the survivor-average number of recurrence (SANR) estimands as
@@
where $\mu^{z}\quantity(t;r)=\mathbb{E}\quantity[N_{i}^{z}(t)\mid\mathcal{AS}(r)]$ for $t<r$. The function $g(\cdot,\cdot)$ determines the scale of the effect measure. For example, g(x,y)=xyg(x,y)=x-y and $g(x,y)=x/y$ correspond to causal mean difference and causal rate ratio, respectively. This class of estimands represents the causal contrast in the number of event occurrences before time $t$ for the subset of patients who would survive until time $r$ under both treatment arms. This estimand generalizes those considered in Lyu et al. (2023), to which it reduces when $r=t$ and $g(x,y)=x/y$. Table 3 summarizes how our definitions of the estimands and the always-survivor stratum differ from those in the existing literature, thereby positioning the contribution of this paper relative to prior work on principal stratification.
Table 3: Summary of literature on principal stratification in continuous time. We summarize the following features: (i) type of outcome, (ii) scale of estimand, (iii) definition of always-survivor stratum, (iv) cross-world assumption (v) whether cross-world and within-world dependence is distinguished, (vi) identification strategy (point identification vs partial identification with sensitivity analysis (SA)), and (vii) inference model.
Nevo and Gorfine (2022) Xu et al. (2022) Lyu et al. (2023) Comment et al. (2025) This article Outcome type Time to event Time to event Recurrent event Time to event Recurrent event Estimand scale Difference Ratio Ratio Difference General scale Always-survivor stratum Population stratification Single index Single index Double index Double index Cross-world assumption Frailty Copula Frailty Frailty Frailty Cross- and within-world Distinguished Distinguished Conflated Conflated Distinguished Identification Partial with SA Partial with SA Point Point Partial with SA Model Parametric Nonparameteric Parametric Parametric Nonparameteric
@@ 3 Bayesian Nonparametric Causal Inference
3.1 Overview
Section 2 defined the potential terminal time $D_{i}^{z}$ and the recurrent-event schedule $\{T_{ij}^{z}\}_{j\geq 1}$ (equivalently $N_{i}^{z}(\cdot)$) under each arm $z\in\{0,1\}$. We work on the log-time scale in our Bayesian model: $U_{i}^{z}=\log(D_{i}^{z})$ and $Y_{ij}^{z}=\log(W_{ij}^{z})$ with $W_{ij}^{z}=T_{ij}^{z}-T_{i(j-1)}^{z}$. If unit $i$ is assigned to arm $Z_{i}=z$, then each observed gap time $W_{ij}=T_{ij}-T_{i(j-1)}$ corresponds to the $j$th gap under arm $z$, so that by Assumption 1 we observe $Y_{ij}=\log(W_{ij})=Y_{ij}^{z}=\log(W_{ij}^{z})$ for $j=1,\ldots,N_{i}$. The next gap under the assigned arm, $Y_{i,N_{i}+1}=Y_{i,N_{i}+1}^{Z_{i}}$, is right-censored at $L_{i,N_{i}+1}=\log(\mathcal{T}_{i}-T_{iN_{i}})$ because no additional recurrence is observed in $(T_{iN_{i}},\mathcal{T}_{i}]$; this censoring is induced either by the terminal event ($\mathcal{T}_{i}=D_{i}^{Z_{i}}$) or by other censoring mechanisms ($\mathcal{T}_{i}=C_{i}^{Z_{i}}$). Likewise, the terminal log-time under the assigned arm, $U_{i}=U_{i}^{Z_{i}}=\log(D_{i}^{Z_{i}})$, is observed when $\delta_{i}^{C}=0$ (death observed) and is right-censored at $\log(\mathcal{T}_{i})$ when $\delta_{i}^{C}=1$. Therefore, the observed-data likelihood (defined later in Section 3.3) is exactly the likelihood for the variables $\{U_{i}^{Z_{i}},Y_{i1}^{Z_{i}},Y_{i2}^{Z_{i}},\ldots\}$ under this right-censoring structure. Because $(U_{i}^{z},\{Y_{ij}^{z}\}_{j\geq 1})$ is a one-to-one transformation of $(D_{i}^{z},\{T_{ij}^{z}\}_{j\geq 1})$, a model for $(U_{i}^{z},\{Y_{ij}^{z}\}_{j\geq 1})$ induces a model for $(D_{i}^{z},N_{i}^{z}(\cdot))$ and thus for the causal estimands.
Let $\mathbf{Y}_{i}^{z}=(Y_{i1}^{z},Y_{i2}^{z},\ldots)$ denote the infinite log-gap sequence under arm $z$. Bayesian inference targets the joint distribution of all variables $(\mathbf{Y}^{0},\mathbf{U}^{0},\mathbf{Y}^{1},\mathbf{U}^{1},\mathbf{C}^{0},\mathbf{C}^{1},\mathbf{Z},\mathbf{X})$. We assume a priori independence between parameters governing $(\mathbf{C}^{0},\mathbf{C}^{1},\mathbf{Z},\mathbf{X})$ and those governing $(\mathbf{Y}^{0},\mathbf{U}^{0},\mathbf{Y}^{1},\mathbf{U}^{1})$. Under Assumption 3 and this prior independence, the censoring mechanism is ignorable for inference about the recurrent/terminal potential outcomes (e.g., Rubin, 1978), so we focus on $(\mathbf{Y}^{0},\mathbf{U}^{0},\mathbf{Y}^{1},\mathbf{U}^{1},\mathbf{Z},\mathbf{X})$.
Assuming unit exchangeability, de Finetti’s theorem implies the existence of global parameters $\boldsymbol{\theta}^{*}$ and $\boldsymbol{\varphi}^{*}$ such that
@@ 3.1 Overview
$\begin{split}&p\!\left(\mathbf{Y}^{0},\mathbf{U}^{0},\mathbf{Y}^{1},\mathbf{U}^{1},\mathbf{Z},\mathbf{X}\right)\\ &\qquad=\int p(\boldsymbol{\theta}^{*})\,p(\boldsymbol{\varphi}^{*})\prod_{i=1}^{n}p\!\left(\mathbf{Y}_{i}^{0},U_{i}^{0},\mathbf{Y}_{i}^{1},U_{i}^{1}\mid\mathbf{X}_{i},\boldsymbol{\theta}^{*}\right)\,p\!\left(Z_{i},\mathbf{X}_{i}\mid\boldsymbol{\varphi}^{*}\right)\,d\boldsymbol{\theta}^{*}\,d\boldsymbol{\varphi}^{*}.\end{split}$ (2)
@@
where the global parameter $\boldsymbol{\theta}^{*}$ has prior $p(\boldsymbol{\theta}^{*})$. Under the ignorable treatment assignment (Assumption 2), the assignment model does not affect the inference of the quantities of scientific interest $\boldsymbol{\theta}^{*}$ either. In what follows, we will conduct a likelihood analysis for $\boldsymbol{\theta}^{*}$, assuming that the value of $\boldsymbol{\theta}^{*}$ which governed the distribution of observable data has been drawn from a prior distribution with compact support. For identification, we further assume that a frailty $\boldsymbol{\gamma}$ is included as unknown parameters in an expanded parameter set $\boldsymbol{\theta}^{*}=(\boldsymbol{\theta},\boldsymbol{\gamma})$ for $\boldsymbol{\gamma}=(\boldsymbol{\gamma}_{1},\ldots,\boldsymbol{\gamma}_{n})^{\top}$, as formalized in the following section.
3.2 Identification via frailty and sensitivity analysis
The causal estimands depend on the joint distribution of the potential outcomes across arms, which is not identified from observed data alone because $(\mathbf{Y}_{i}^{0},U_{i}^{0})$ and $(\mathbf{Y}_{i}^{1},U_{i}^{1})$ are never jointly observed. To resolve this identification challenge, we introduce a bivariate frailty $\boldsymbol{\gamma}_{i}=(\gamma_{i}^{0},\gamma_{i}^{1})$ to transparently parameterize cross-world dependence.
Assumption 4 (Cross-world dependence via bivariate frailty).
@@ 3.2 Identification via frailty and sensitivity analysis
For each unit $i$, there exists a bivariate frailty $\boldsymbol{\gamma}_{i}=(\gamma_{i}^{0},\gamma_{i}^{1})$ such that
$p\!\left(\mathbf{Y}_{i}^{0},U_{i}^{0},\mathbf{Y}_{i}^{1},U_{i}^{1}\mid\mathbf{X}_{i},\boldsymbol{\gamma}_{i}\right)$ $=p\!\left(\mathbf{Y}_{i}^{0},U_{i}^{0}\mid\mathbf{X}_{i},\gamma_{i}^{0}\right)\;p\!\left(\mathbf{Y}_{i}^{1},U_{i}^{1}\mid\mathbf{X}_{i},\gamma_{i}^{1}\right)\text{ and }$
$p\!\left(\mathbf{Y}_{i}^{z},U_{i}^{z}\mid\mathbf{X}_{i},\gamma_{i}^{z}\right)$ $=p\!\left(\mathbf{Y}_{i}^{z}\mid\mathbf{X}_{i},\gamma_{i}^{z}\right)p\!\left(U_{i}^{z}\mid\mathbf{X}_{i},\gamma_{i}^{z}\right)\text{ for }z=0,1.$
Although the frailty device has been previously adopted in the semi-competing risks literature (e.g., Stensrud et al., 2017; Nevo and Gorfine, 2022; Lyu et al., 2023; Comment et al., 2025), our approach differs from the existing literature in several important respects. Table 3 summarizes the key distinctions between our frailty-based approach and prior work.
Under Assumption 4, the individual frailty vector $\boldsymbol{\gamma}_{i}=(\gamma_{i}^{0},\gamma_{i}^{1})$ governs two distinct forms of dependence: (i) cross-world dependence between treatment arms, arising from the association between $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$; and (ii) within-world dependence between the terminal event $U_{i}^{z}$ and the recurrent event process $\mathbf{Y}_{i}^{z}$ under treatment $z$. Component (i) is unidentifiable because $(\mathbf{Y}_{i}^{0},U_{i}^{0})$ and $(\mathbf{Y}_{i}^{1},U_{i}^{1})$ are never jointly observable, whereas Component (ii), which captures within-world dependence between recurrent events and the terminal event, has been recognized as an important source of variation in noncausal, associative analyses, where cross-world dependence is not considered (e.g., Paulon et al., 2020; Tian et al., 2024).
@@
A common simplification sets $\gamma_{i}^{0}=\gamma_{i}^{1}\equiv\gamma_{i}$ (e.g., Lyu et al., 2023; Comment et al., 2025), thereby assuming cross-world independence in Component (i) and yielding point identification of the causal contrasts. In causal analyses, however, this restriction also imposes a common within-world dependence structure between recurrent events and the terminal event across treatment worlds, an assumption that is unlikely to hold unless all effect modifiers are correctly measured and modeled. Therefore, this approach conflates two distinct dependence structures, one inherently non-identifiable (Component (i)) and the other identifiable (Component (ii)), thereby imposing a strong and opaque assumption in causal settings.
@@
Our approach retains separate frailties $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$ and introduces a sensitivity parameter $\rho=\operatorname{Cor}(\gamma_{i}^{0},\gamma_{i}^{1})$ that captures their cross-world association, similar to Nevo and Gorfine (2022) who addressed the non-recurrent event setting. This formulation cleanly separates identifiable and non-identifiable components: the marginal distributions of $\gamma_{i}^{0}$ and $\gamma_{i}^{1}$, and hence the within-world dependence, are estimable from the observed data, whereas the cross-world correlation $\rho$ is an unidentifiable quantity. Nevo and Gorfine (2022) considered a non-recurrent event setting and specified a correlated gamma distribution for the joint frailty, fixing the location parameters and correlation parameter $\rho$ a priori and thereby constraining the model to a relatively restrictive parametric family. In contrast, our method places a flexible nonparametric prior on the joint distribution of $(\gamma_{i}^{0},\gamma_{i}^{1})$ conditional on a fixed $\rho$, specified in Section 3.3.2.
Given $\rho$, posterior inference proceeds by sampling $(\boldsymbol{\theta},\boldsymbol{\gamma})$ from $p(\boldsymbol{\theta},\boldsymbol{\gamma}\mid\mathbf{O})\propto p(\boldsymbol{\theta},\boldsymbol{\gamma}\mid\rho)\prod_{i=1}^{n}p\!\left(U_{i},\mathbf{Y}_{i}\mid Z_{i},\mathbf{X}_{i},\boldsymbol{\theta},\gamma_{i}^{Z_{i}}\right)$ and imputing missing potential outcomes $(U_{i}^{1-Z_{i}},\mathbf{Y}_{i}^{1-Z_{i}})$ from their posterior predictive distribution. This yields posterior draws of the always-survivor indicators and the causal estimands defined in Section 2.2. Although we denote $\mathbf{Y}_{i}^{z}=(Y_{i1}^{z},Y_{i2}^{z},\ldots)$ as an infinite sequence, computation only requires simulating the latent schedule up to the largest horizon $r$ used to define the principal strata and estimands (and for $t\leq r$). Theorem 1 states that (1) is identifiable up to the frailty $\boldsymbol{\gamma}$.